# Existence of variational solutions to doubly nonlinear systems in nondecreasing domains

### v1

@@ 4.6. Variational inequality for the limit map
$+\iint_{\Omega\times(mh_{\ell},\tau]}f(x,u_{\ell,m},Du_{\ell,m})\,\mathrm{d}x\mathrm{d}t$
$\leq\sum_{i=1}^{m}\iint_{\Omega\times(t_{\ell,i-1},t_{\ell,i}]}f(x,w,Dw)+\tfrac{1}{h_{\ell}}\big{(}\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\cdot(w-u_{\ell,i})\,\mathrm{d}x\mathrm{d}t$
$\phantom{=}+\iint_{\Omega\times(mh_{\ell},\tau]}f(x,w,Dw)+\tfrac{1}{h_{\ell}}\big{(}\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\cdot(w-u_{\ell,i})\,\mathrm{d}x\mathrm{d}t$
(4.12) $=\iint_{\Omega_{\tau}}f(x,w,Dw)+\Delta_{-h_{\ell}}\big{(}\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\cdot(w-u_{\ell})\,\mathrm{d}x\mathrm{d}t$
@@
(4.13) $\leq\iint_{\Omega_{\tau}}f(x,v,Dv)\,\mathrm{d}x\mathrm{d}t+\iint_{\Omega_{\tau}}\Delta_{h_{\ell}}v\cdot\big{(}\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t$
@@
$\delta_{2}(h_{\ell})$ $:=\iint_{\Omega\times(-h_{\ell},0)}\Delta_{h_{\ell}}v\cdot\big{(}\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}(t+h_{\ell})-\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$q$}}}}(t)\big{)}\,\mathrm{d}x\mathrm{d}t.$
@@
[…]
@@
Next, note that $\Delta_{h_{\ell}}v\to\partial_{t}v$ in $L^{q+1}(\Omega_{T},\mathds{R}^{N})$. Moreover, since $\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$q$}}}}$ is bounded in $L^{\frac{q+1}{q}}(\Omega_{T},\mathds{R}^{N})$ by (4.4), and since $\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$q$}}}}\to\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}$ a.e. in $\Omega_{T}$ as $\mathfrak{K}\ni\ell\to\infty$ by (4.11), there exists a (not relabeled) subsequence such that $\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$q$}}}}\rightharpoondown\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}$ as $\mathfrak{K}\ni\ell\to\infty$. This gives us that
(4.16) $\lim_{\mathfrak{K}\ni\ell\to\infty}\iint_{\Omega_{\tau}}\Delta_{h_{\ell}}v\cdot\big{(}\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t=\iint_{\Omega_{\tau}}\partial_{t}v\cdot\big{(}\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t$
@@
[…]
@@
(4.19) $\leq\iint_{\Omega_{t_{o}+\delta}}f(x,v,Dv)\,\mathrm{d}x\mathrm{d}t+\mathchoice{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}\!\int_{t_{o}}^{t_{o}+\delta}\iint_{\Omega_{\tau}}\partial_{t}v\cdot\big{(}\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t\,\mathrm{d}\tau$
[…]
@@
$\iint_{\Omega_{t_{o}}}f(x,u,Du)\,\mathrm{d}x\mathrm{d}t\leq$ $\iint_{\Omega_{t_{o}}}f(x,v,Dv)\,\mathrm{d}x\mathrm{d}t\mathrm{d}\tau+\iint_{\Omega_{t_{o}}}\partial_{t}v\cdot\big{(}\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t$
@@ 5.1. Preparation
Assume that $E\subset\Omega\times[0,T)$ is a relatively open noncylindrical domain, and let $f\colon\Omega\times\mathds{R}^{N}\times\mathds{R}^{Nn}\to[0,\infty]$ be a Carathéodory function. We say that $u\in V^{p}(E)$ with $\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\in L^{1}(E,\mathds{R}^{N})$ is a parabolic minimizer if and only if
$\iint_{E}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot\partial_{t}\phi+f(x,u,Du)\,\mathrm{d}x\mathrm{d}t\leq\iint_{E}f(x,u+\phi,Du+D\phi)\,\mathrm{d}x\mathrm{d}t$
@@
holds by (2.10). Further, since $E$ is nondecreasing and $\operatorname{spt}\big{(}u_{o}^{(\varepsilon)}-u_{\ast}\big{)}\subset E^{0}$ by construction, it follows that $\operatorname{spt}\big{(}u_{o}^{(\varepsilon)}-u_{\ast}\big{)}\subset E^{t}$ for all $t\in[0,T]$. Moreover, we define $[u]_{h}(t)$ according to (3.7) with initial values $u_{o}^{(\varepsilon)}$, and for a test function $\phi\in C_{0}^{\infty}(E,\mathds{R}^{N})$, let $[\phi]_{h}(t)$ be defined according to (3.7) with zero initial values. Due to (2.2), for $v_{h}(t):=[u]_{h}(t)+s[\phi]_{h}(t)$ with $s>0$ we find that $v_{h}=u_{\ast}$ a.e. in $\Omega_{T}\setminus E$. Hence, $v_{h}$ is an admissible comparison map in the variational inequality (2.8), which holds for any $\tau\in[0,T]$ and thus in particular for $\tau=T$, since $u\in C^{0}([0,T];L^{q+1}(\Omega,\mathds{R}^{N}))$ by Proposition 3.16. Omitting the non-negative boundary term on the left-hand side, we obtain that
@@
(5.1) $\leq\iint_{E}f(x,v_{h},Dv_{h})\,\mathrm{d}x\mathrm{d}t+\iint_{E}\partial_{t}v_{h}\cdot\big{(}\boldsymbol{v_{h}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t+\int_{E^{0}}\mathfrak{b}\big{[}u_{o},u_{o}^{(\varepsilon)}\big{]}\,\mathrm{d}x.$
@@
$\iint_{E}$ $\partial_{t}v_{h}\cdot\big{(}\boldsymbol{v_{h}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t$
$=\iint_{\Omega_{T}}\partial_{t}v_{h}\cdot\boldsymbol{v_{h}^{\mbox{\unboldmath{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t-\iint_{\Omega_{T}}\partial_{t}[u]_{h}\cdot\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t-s\iint_{E}\partial_{t}[\phi]_{h}\cdot\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t$
@@
Let $\psi(\cdot):=\tfrac{1}{q+1}|\cdot|^{q+1}$. Using $\psi^{\prime}(v_{h})=\boldsymbol{v_{h}^{\mbox{\unboldmath{\scriptsize$q$}}}}$, we compute
@@
$-\partial_{t}[u]_{h}\cdot\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}$ $=\tfrac{1}{h}([u]_{h}-u)\psi^{\prime}(u)\leq\tfrac{1}{h}(\psi([u]_{h})-\psi(u))$
@@
$\iint_{E}$ $\partial_{t}v_{h}\cdot\big{(}\boldsymbol{v_{h}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t$
$\leq\int_{\Omega\times\{T\}}\psi([u]_{h}+s[\phi]_{h})-[\psi(u)]_{h}\,\mathrm{d}x-s\iint_{E}\partial_{t}[\phi]_{h}\cdot\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t.$
@@
$\limsup_{h\downarrow 0}\iint_{E}\partial_{t}v_{h}\cdot(\boldsymbol{v_{h}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}})\,\mathrm{d}x\mathrm{d}t\leq-s\iint_{E}\partial_{t}\phi\cdot\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t.$
@@
$\leq\iint_{E}f(x,u+s\phi,Du+sD\phi)\,\mathrm{d}x\mathrm{d}t-s\iint_{E}\partial_{t}\phi\cdot\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t+\int_{E^{0}}\mathfrak{b}\big{[}u_{o},u_{o}^{(\varepsilon)}\big{]}\,\mathrm{d}x$
$\leq\iint_{E}(1-s)f(x,u,Du)+sf(x,u+\phi,Du+D\phi)\,\mathrm{d}x\mathrm{d}t-s\iint_{E}\partial_{t}\phi\cdot\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t$
@@
$\iint_{E}\partial_{t}\phi\cdot\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}+f(x,u,Du)\,\mathrm{d}x\mathrm{d}t\leq\iint_{E}f(x,u+\phi,Du+D\phi)\,\mathrm{d}x\mathrm{d}t.$
@@
$\iint_{E}$ $\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot\partial_{t}\phi\,\mathrm{d}x\mathrm{d}t$
@@
in the limit $s\downarrow 0$, for any $\phi\in C_{0}^{\infty}(E,\mathds{R}^{N})$. Since we can replace $\phi$ by $-\phi$ above, we obtain the same estimate for the absolute value of the left-hand side integral. Therefore, applying Hölder’s inequality leads to
$\bigg{|}\iint_{E}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot\partial_{t}\phi\,\mathrm{d}x\mathrm{d}t\bigg{|}$ $\leq c\iint_{E}\big{(}(|Du|+|u|)^{p-1}+|G|^{\frac{p-1}{p}}\big{)}(|D\phi|+|\phi|)\,\mathrm{d}x\mathrm{d}t$
@@
for all $\phi\in C_{0}^{\infty}(E,\mathds{R}^{N})$. Since (2.2) and (2.9) hold, by Remark 3.21 we know that $C_{0}^{\infty}(E,\mathds{R}^{N})$ is dense in $V^{p,0}(E)$. Therefore, we obtain that $\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\in\big{(}V^{p,0}(E)\big{)}^{\prime}$, as well as the claimed estimate in the $\big{(}V^{p,0}(E)\big{)}^{\prime}$-norm.

### v2

@@ 4.6. Variational inequality for the limit map
[…]
$\leq\sum_{i=1}^{m}\iint_{\Omega\times(t_{\ell,i-1},t_{\ell,i}]}f(x,w,Dw)+\tfrac{1}{h_{\ell}}\big(\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$q$}}}}\big)\cdot(w-u_{\ell,i})\,\mathrm{d}x\mathrm{d}t$
$\phantom{=}+\iint_{\Omega\times(mh_{\ell},\tau]}f(x,w,Dw)+\tfrac{1}{h_{\ell}}\big(\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$q$}}}}\big)\cdot(w-u_{\ell,i})\,\mathrm{d}x\mathrm{d}t$
(4.12) $=\iint_{\Omega_{\tau}}f(x,w,Dw)+\Delta_{-h_{\ell}}\big(\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$q$}}}}\big)\cdot(w-u_{\ell})\,\mathrm{d}x\mathrm{d}t$
@@
(4.13) $\leq\iint_{\Omega_{\tau}}f(x,v,Dv)\,\mathrm{d}x\mathrm{d}t+\iint_{\Omega_{\tau}}\Delta_{h_{\ell}}v\cdot\big(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t$
@@
$\delta_{2}(h_{\ell})$ $:=\iint_{\Omega\times(-h_{\ell},0)}\Delta_{h_{\ell}}v\cdot\big(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}(t+h_{\ell})-\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$q$}}}}(t)\big)\,\mathrm{d}x\mathrm{d}t.$
@@
$\phantom{=}+\mathchoice{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}\!\int_{t_{o}}^{t_{o}+\delta}\iint_{\Omega_{\tau}}\Delta_{h_{\ell}}v\cdot\big(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t\,\mathrm{d}\tau$
@@
Next, note that $\Delta_{h_{\ell}}v\to\partial_{t}v$ in $L^{q+1}(\Omega_{T},\mathds{R}^{N})$. Moreover, since $\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$q$}}}}$ is bounded in $L^{\frac{q+1}{q}}(\Omega_{T},\mathds{R}^{N})$ by (4.4), and since $\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$q$}}}}\to\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}$ a.e. in $\Omega_{T}$ as $\mathfrak{K}\ni\ell\to\infty$ by (4.11), there exists a (not relabeled) subsequence such that $\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$q$}}}}\rightharpoondown\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}$ as $\mathfrak{K}\ni\ell\to\infty$. This gives us that
(4.16) $\lim_{\mathfrak{K}\ni\ell\to\infty}\iint_{\Omega_{\tau}}\Delta_{h_{\ell}}v\cdot\big(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t=\iint_{\Omega_{\tau}}\partial_{t}v\cdot\big(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t$
@@
(4.17) $\mathchoice{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}\!\int_{t_{o}}^{t_{o}+\delta}\int_{\Omega}\mathfrak{b}[u(t),v(t)]\,\mathrm{d}x\mathrm{d}t\leq\liminf_{\mathfrak{K}\ni\ell\to\infty}\tfrac{1}{\delta}\int_{t_{o}}^{t_{o}+\delta-h_{\ell}}\int_{\Omega}\mathfrak{b}[u_{\ell}(t),v(t+h_{\ell})]\,\mathrm{d}x\mathrm{d}t.$
@@
(4.19) $\leq\iint_{\Omega_{t_{o}+\delta}}f(x,v,Dv)\,\mathrm{d}x\mathrm{d}t+\mathchoice{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}\!\int_{t_{o}}^{t_{o}+\delta}\iint_{\Omega_{\tau}}\partial_{t}v\cdot\big(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t\,\mathrm{d}\tau$
$\phantom{=}-\mathchoice{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}\!\int_{t_{o}}^{t_{o}+\delta}\int_{\Omega}\mathfrak{b}[u(t),v(t)]\,\mathrm{d}x\mathrm{d}t\,\mathrm{d}\tau+\int_{\Omega}\mathfrak{b}[u_{o},v(0)]\,\mathrm{d}x$
@@
$\iint_{\Omega_{t_{o}}}f(x,u,Du)\,\mathrm{d}x\mathrm{d}t\leq$ $\iint_{\Omega_{t_{o}}}f(x,v,Dv)\,\mathrm{d}x\mathrm{d}t\mathrm{d}\tau+\iint_{\Omega_{t_{o}}}\partial_{t}v\cdot\big(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t$
@@ 5.1. Preparation
Assume that $E\subset\Omega\times[0,T)$ is a relatively open noncylindrical domain, and let $f\colon\Omega\times\mathds{R}^{N}\times\mathds{R}^{Nn}\to[0,\infty]$ be a Carathéodory function. We say that $u\in V^{p}(E)$ with $\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\in L^{1}(E,\mathds{R}^{N})$ is a parabolic minimizer if and only if
$\iint_{E}\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\cdot\partial_{t}\phi+f(x,u,Du)\,\mathrm{d}x\mathrm{d}t\leq\iint_{E}f(x,u+\phi,Du+D\phi)\,\mathrm{d}x\mathrm{d}t$
@@
holds by (2.10). Further, since $E$ is nondecreasing and $\operatorname{spt}\big(u_{o}^{(\varepsilon)}-u_{\ast}\big)\subset E^{0}$ by construction, it follows that $\operatorname{spt}\big(u_{o}^{(\varepsilon)}-u_{\ast}\big)\subset E^{t}$ for all $t\in[0,T]$. Moreover, we define $u]_{h}(t)$ according to (3.7) with initial values $u_{o}^{(\varepsilon)}$, and for a test function $\phi\in C_{0}^{\infty}(E,\mathds{R}^{N})$, let $\phi]_{h}(t)$ be defined according to (3.7) with zero initial values. Due to (2.2), for $v_{h}(t):=[u]_{h}(t)+s[\phi]_{h}(t)$ with $s>0$ we find that $v_{h}=u_{\ast}$ a.e. in $\Omega_{T}\setminus E$. Hence, $v_{h}$ is an admissible comparison map in the variational inequality (2.8), which holds for any $\tau\in[0,T]$ and thus in particular for $\tau=T$, since $u\in C^{0}([0,T];L^{q+1}(\Omega,\mathds{R}^{N}))$ by Proposition 3.16. Omitting the non-negative boundary term on the left-hand side, we obtain that
@@
(5.1) $\leq\iint_{E}f(x,v_{h},Dv_{h})\,\mathrm{d}x\mathrm{d}t+\iint_{E}\partial_{t}v_{h}\cdot\big(\boldsymbol{v_{h}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t+\int_{E^{0}}\mathfrak{b}\big[u_{o},u_{o}^{(\varepsilon)}\big]\,\mathrm{d}x.$
@@
$\iint_{E}$ $\partial_{t}v_{h}\cdot\big(\boldsymbol{v_{h}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t$
$=\iint_{\Omega_{T}}\partial_{t}v_{h}\cdot\boldsymbol{v_{h}^{\mbox{{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t-\iint_{\Omega_{T}}\partial_{t}[u]_{h}\cdot\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t-s\iint_{E}\partial_{t}[\phi]_{h}\cdot\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t$
@@
Let $\psi(\cdot):=\tfrac{1}{q+1}|\cdot|^{q+1}$. Using $\psi^{\prime}(v_{h})=\boldsymbol{v_{h}^{\mbox{{\scriptsize$q$}}}}$, we compute
@@
$-\partial_{t}[u]_{h}\cdot\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}$ $=\tfrac{1}{h}([u]_{h}-u)\psi^{\prime}(u)\leq\tfrac{1}{h}(\psi([u]_{h})-\psi(u))$
@@
$\iint_{E}$ $\partial_{t}v_{h}\cdot\big(\boldsymbol{v_{h}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t$
$\leq\int_{\Omega\times\{T\}}\psi([u]_{h}+s[\phi]_{h})-[\psi(u)]_{h}\,\mathrm{d}x-s\iint_{E}\partial_{t}[\phi]_{h}\cdot\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t.$
@@
$\limsup_{h\downarrow 0}\iint_{E}\partial_{t}v_{h}\cdot(\boldsymbol{v_{h}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}})\,\mathrm{d}x\mathrm{d}t\leq-s\iint_{E}\partial_{t}\phi\cdot\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t.$
@@
$\leq\iint_{E}f(x,u+s\phi,Du+sD\phi)\,\mathrm{d}x\mathrm{d}t-s\iint_{E}\partial_{t}\phi\cdot\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t+\int_{E^{0}}\mathfrak{b}\big[u_{o},u_{o}^{(\varepsilon)}\big]\,\mathrm{d}x$
$\leq\iint_{E}(1-s)f(x,u,Du)+sf(x,u+\phi,Du+D\phi)\,\mathrm{d}x\mathrm{d}t-s\iint_{E}\partial_{t}\phi\cdot\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\,\mathrm{d}x\mathrm{d}t$
@@
$\iint_{E}\partial_{t}\phi\cdot\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}+f(x,u,Du)\,\mathrm{d}x\mathrm{d}t\leq\iint_{E}f(x,u+\phi,Du+D\phi)\,\mathrm{d}x\mathrm{d}t.$
@@
$\iint_{E}$ $\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\cdot\partial_{t}\phi\,\mathrm{d}x\mathrm{d}t$
@@
in the limit $s\downarrow 0$, for any $\phi\in C_{0}^{\infty}(E,\mathds{R}^{N})$. Since we can replace $\phi$ by $\phi$ above, we obtain the same estimate for the absolute value of the left-hand side integral. Therefore, applying Hölder’s inequality leads to
$\bigg|\iint_{E}\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\cdot\partial_{t}\phi\,\mathrm{d}x\mathrm{d}t\bigg|$ $\leq c\iint_{E}\big((|Du|+|u|)^{p-1}+|G|^{\frac{p-1}{p}}\big)(|D\phi|+|\phi|)\,\mathrm{d}x\mathrm{d}t$
@@
for all $\phi\in C_{0}^{\infty}(E,\mathds{R}^{N})$. Since (2.2) and (2.9) hold, by Remark 3.21 we know that $C_{0}^{\infty}(E,\mathds{R}^{N})$ is dense in $V^{p,0}(E)$. Therefore, we obtain that $\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\in\big(V^{p,0}(E)\big)^{\prime}$, as well as the claimed estimate in the $\big(V^{p,0}(E)\big)^{\prime}$-norm.
@@
Data availability. Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Conflict of Interests. The authors do not have a conflict of interests.